# Possible explanations of the observed $\Lambda_{c}$ resonances

### v1

@@ Abstract
Inspired by the latest experimental progress, we systematically study the OZI-allowed two-body strong decay properties of $1P$-, $1D$-, $2S$- and $2P$-wave $\Lambda_{c}$ baryons within the $j$-$j$ coupling scheme in the framework of the quark pair creation model. The calculations indicate that: ($\rm{i}$) Taking the observed states $\Lambda_{c}(2595)^{+}$ and $\Lambda_{c}(2625)^{+}$ as the $1P$-wave $\lambda$-modes states $\Lambda_{c}|J^{P}=1/2^{-},1\rangle_{\lambda}$ and $\Lambda_{c}|J^{P}=3/2^{-},1\rangle_{\lambda}$, respectively, we can reproduce the experimental data well in theory. ($\rm{ii}$) Combining with the measured mass and the decay properties of $\Lambda_{c}(2860)^{+}$, this excited state can be explained as $1D$-wave $\lambda$-mode state $\Lambda_{c}|J^{P}=3/2^{+},1\rangle_{\lambda\lambda}$. ($\rm{iii}$) The newly observed state $\Lambda_{c}(2910)^{+}$ may be assigned as one of the $1P$-wave $\rho$-mode states $\Lambda_{c}|J^{P}=3/2^{-},2\rangle_{\rho}$ or $\Lambda_{c}|J^{P}=5/2^{-},2\rangle_{\rho}$. Meanwhile, we notice that the partial decay width ratio between $\Sigma_{c}\pi$ and $\Sigma_{c}^{*}\pi$ for the two candidates is significantly different. Hence, experimental progress in this ratio measurement can shed light on the nature of $\Lambda_{c}(2910)^{+}$. ($\rm{iv}$) According to the properties of $\Lambda_{c}(2765)^{+}$, we find that the $2S$-wave $\lambda$-mode state $\Lambda_{c1}|J^{P}=1/2^{+},0\rangle_{\lambda}$ may be a potential candidate. ($\rm{v}$) The $2P$-wave $\lambda$-mode state $\Lambda_{c1}|J^{P}=3/2^{-},1\rangle_{\lambda}$ is mostly likely to be a good assignment of the controversial state $\Lambda_{c}(2940)^{+}$. Both the total decay width and partial decay ratio between $pD^{0}$ and $\Sigma_{c}\pi$ are in good agreement with the observations. ($\rm{vi}$) In addition, for the missing $\Lambda_{c}$ excitations, we obtain their strong decay properties and hope that’s useful for future experimental exploration.
@@ I Introduction
Meanwhile, there exist many theoretical calculations to decode the inner structures of those observed $\Lambda_{c}^{+}$ baryons via mass spectrum [18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32] and decay properties [33, 34, 35, 36, 37, 38, 19, 39, 40, 41, 42, 43, 44, 45]. Except for the $\Lambda_{c}^{+}$ ground state, the $\Lambda_{c}(2595)^{+}$ and $\Lambda_{c}(2625)^{+}$ can be well interpreted as the $P$-wave $\Lambda_{c}$ states of $J^{P}=1/2^{-}$ and $J^{P}=3/2^{-}$ [18, 19, 20, 33, 34, 35, 36, 37, 38], respectively. Meanwhile, for the $\Lambda_{c}(2860)^{+}$ and $\Lambda_{c}(2880)^{+}$, most of the references suggest they together form the $D$-wave $\Lambda_{c}$ doublet of $J^{P}=3/2^{+}$ and $J^{P}=5/2^{+}$, respectively [25, 23, 26, 27, 40, 41]. However, there is some literature that don’t support $\Lambda_{c}(2880)^{+}$ as $J^{P}=5/2^{+}$ state [27, 40], and believe this state being $F$-wave $\Lambda_{c}$ states of $J^{P}=5/2^{-}$ [40]. Hence, the relation between those two states still needs to be carefully examined in future experimental and theoretical studies. As for $\Lambda_{c}(2940)^{+}$, its internal structure is controversial. In addition to being interpreted as a traditional hadron state of $J^{P}=1/2^{-}$ [2], $3/2^{\pm}$ [22, 24, 40, 42], $5/2^{\mp}$ [42, 44] or $7/2^{+}$ [43], it has also been interpreted as $D^{*}N$ molecular state with $J^{P}=3/2^{-}$ [46, 47]. Fortunately, the latest experimental measurements by Belle Collaboration [16] about the partial decay width ratios of $\Lambda_{c}(2940)^{+}$ will provide a stronger basis for decoding its inner structure. Compared to $\Lambda_{c}(2940)^{+}$, the properties of $\Lambda_{c}(2765)^{+}$(or $\Sigma_{c}(2765)^{+}$) and $\Lambda_{c}(2910)^{+}$ are more controversial, and it is not even certain whether they are $\Lambda_{c}$ or $\Sigma_{c}$ states. At present, theoretical explanations suggest that $\Lambda_{c}(2765)^{+}$(or $\Sigma_{c}(2765)^{+}$) may be $\Lambda_{c}(2S)1/2^{+}$ [21, 22], $\Lambda_{c}(1P)1/2^{-}$ [22] or $\Sigma_{c}(1P)3/2^{-}$ [21] resonance. For $\Lambda_{c}(2910)^{+}$, it can be assigned as $\Lambda_{c}$ resonance with spin-parity $1/2^{+}$ [22] or $1/2^{-}$ [24]. Moreover, the assignment of the two states as $D^{(*)}N$ molecular states also exists [47, 48].
Table 1: Predicted masses of $\Lambda_{c}$ states($1P$-, $1D$-, $2S$- and $2P$-wave) in various quark models and possible decay channels within the quark pair creation model.
@@
$\Lambda_{c}$|${\ket{J^{P},j}}_{\lambda(\rho)}$ $n_{\lambda}$ $n_{\rho}$ $l_{\lambda}$ $l_{\rho}$ $L$ $s_{\rho}$ $j$ $J^{P}$ RQM [28] RQM [30] RFT [31] NCQM [25] QM [32]
$\Lambda_{c}$|${\ket{J^{P}=\frac{1}{2}^{-},1}}_{\lambda}$ 0 0 1 0 1 0 1 $\frac{1}{2}^{-}$ $2598$ $2630$ $2591$ 2614 $2625$ $\Sigma_{c}\pi$
$\Lambda_{c}$|${\ket{J^{P}=\frac{3}{2}^{-},1}}_{\lambda}$ 0 0 1 0 1 0 1 $\frac{3}{2}^{-}$ 2627 2640 2629 2639 2636
$\Lambda_{c}$|${\ket{J^{P}=\frac{1}{2}^{-},0}}_{\rho}$ 0 0 0 1 1 1 0 $\frac{1}{2}^{-}$ 2780 2816 $\Sigma^{(*)}_{c}\pi$
$\Lambda_{c}$|${\ket{J^{P}=\frac{1}{2}^{-},1}}_{\rho}$ 0 0 0 1 1 1 1 $\frac{1}{2}^{-}$ 2830 2816
$\Lambda_{c}$|${\ket{J^{P}=\frac{3}{2}^{-},1}}_{\rho}$ 0 0 0 1 1 1 1 $\frac{3}{2}^{-}$ 2840 2830
$\Lambda_{c}$|${\ket{J^{P}=\frac{3}{2}^{-},2}}_{\rho}$ 0 0 0 1 1 1 2 $\frac{3}{2}^{-}$ 2885 2830
$\Lambda_{c}$|${\ket{J^{P}=\frac{5}{2}^{-},2}}_{\rho}$ 0 0 0 1 1 1 2 $\frac{5}{2}^{-}$ 2900 2872
$\Lambda_{c}$|${\ket{J^{P}=\frac{3}{2}^{+},2}}_{\lambda\lambda}$ 0 0 2 0 2 0 2 $\frac{3}{2}^{+}$ $2874$ $2910$ $2857$ 2843 $2887$ $\Sigma^{(*)}_{c}\pi$,$DN$
$\Lambda_{c}$|${\ket{J^{P}=\frac{5}{2}^{+},2}}_{\lambda\lambda}$ 0 0 2 0 2 0 2 $\frac{5}{2}^{+}$ 2880 2910 2879 2851 2887
$\Lambda_{c}$|${\ket{J^{P}=\frac{3}{2}^{+},2}}_{\rho\rho}$ 0 0 0 2 2 0 2 $\frac{3}{2}^{+}$ $3035$ $3073$ $\Lambda^{+}_{c}\omega^{0}$,$\Sigma^{(*)}_{c}\pi$,$\Xi^{{}^{\prime}}_{c}K$,$\Sigma_{c}(1P_{\lambda})\pi$
$\Lambda_{c}$|${\ket{J^{P}=\frac{5}{2}^{+},2}}_{\rho\rho}$ 0 0 0 2 2 0 2 $\frac{5}{2}^{+}$ 3140 3092
$\Lambda_{c1}$|${\ket{J^{P}=\frac{1}{2}^{+},0}}_{\lambda}$ 1 0 0 0 0 0 0 $\frac{1}{2}^{+}$ $2769$ $2775$ $2766$ 2772 $2791$ $\Sigma^{(*)}_{c}\pi$,$DN$
$\Lambda_{c1}$|${\ket{J^{P}=\frac{1}{2}^{+},0}}_{\rho}$ 0 1 0 0 0 0 0 $\frac{1}{2}^{+}$ $2970$ $\Sigma^{(*)}_{c}\pi$,$D^{(*)}N$,$\Sigma_{c}(1P_{\lambda})\pi$
$\Lambda_{c1}$|${\ket{J^{P}=\frac{1}{2}^{-},1}}_{\lambda}$ 1 0 1 0 1 0 1 $\frac{1}{2}^{-}$ $2983$ $3030$ $2989$ 2980 $\Sigma^{(*)}_{c}\pi$,$D^{(*)}N$,$\Sigma_{c}(1P_{\lambda})\pi$
$\Lambda_{c1}$|${\ket{J^{P}=\frac{3}{2}^{-},1}}_{\lambda}$ 1 0 1 0 1 0 1 $\frac{3}{2}^{-}$ $3005$ $3035$ $3000$ 3004
$\Lambda_{c1}$|${\ket{J^{P}=\frac{1}{2}^{-},0}}_{\rho}$ 0 1 0 1 1 1 0 $\frac{1}{2}^{-}$ $3200$ $\Lambda^{+}_{c}\omega^{0}/\eta$,$\Sigma^{(*)}_{c}\pi$,$\Xi_{c}K$,$\Xi_{c}^{\prime(*)}K$,$\Sigma_{c}(1P_{\lambda/\rho})\pi$, $\Lambda^{+}_{c}(1P_{\lambda})\eta$
$\Lambda_{c1}$|${\ket{J^{P}=\frac{1}{2}^{-},1}}_{\rho}$ 0 1 0 1 1 1 1 $\frac{1}{2}^{-}$ $3240$
$\Lambda_{c1}$|${\ket{J^{P}=\frac{3}{2}^{-},1}}_{\rho}$ 0 1 0 1 1 1 1 $\frac{3}{2}^{-}$ $3240$
$\Lambda_{c1}$|${\ket{J^{P}=\frac{3}{2}^{-},2}}_{\rho}$ 0 1 0 1 1 1 2 $\frac{3}{2}^{-}$ $3255$
$\Lambda_{c1}$|${\ket{J^{P}=\frac{5}{2}^{-},2}}_{\rho}$ 0 1 0 1 1 1 2 $\frac{5}{2}^{-}$ $3130$
@@ II Theoretical framework
The quark pair creation model [49, 50, 51, 52, 53] as a phenomenological method has been employed successfully in the description of the Okubo-Zweig-Iizuka-allowed two-body strong decays. The main idea of this model is that the quark-antiquark pair with $0^{++}$ is created from the vacuum and then regroups with the quarks from the initial hadron to produce two outing hadrons. Hence, for the $\Lambda_{c}$ system, there are three decay processes, as shown in Fig 1.
@@
$T=$ $-3\gamma\sum\limits_{m}\langle 1m;1-m|00\rangle\int d^{3}\textbf{p}_{4}d^{3}\textbf{p}_{5}\delta^{3}(\textbf{p}_{4}+\textbf{p}_{5})$ (1)
×$\times\omega^{45}_{0}\chi^{45}_{1,-m}\phi^{45}_{0}\mathcal{Y}^{m}_{1}(\dfrac{\textbf{p}_{4}-\textbf{p}_{5}}{2})a^{\mathcal{y}}_{4i}(\textbf{p}_{4})b^{\mathcal{y}}_{5j}(\textbf{p}_{5}).$
@@
×$\times\int d^{3}\mathbf{p}_{1}d^{3}\mathbf{p}_{2}d^{3}\mathbf{p}_{3}\delta^{3}(\mathbf{p}_{1}+\mathbf{p}_{2}+\mathbf{p}_{3}-\mathbf{p}_{A})$
×$\times\Psi_{N_{A}L_{A}M_{L_{A}}(\mathbf{p}_{1},\mathbf{p}_{2},\mathbf{p}_{3})}\chi^{123}_{S_{A}M_{S_{A}}}|q_{1}(\mathbf{p}_{1})q_{2}(\mathbf{p}_{2})q_{3}(\mathbf{p}_{3})\rangle,$ (2)
@@
×$\times\int d^{3}\mathbf{p}_{a}d^{3}\mathbf{p}_{b}\delta^{3}(\mathbf{p}_{a}+\mathbf{p}_{b}-\mathbf{p}_{C})$
×$\times\Psi_{N_{C}L_{C}M_{L_{C}}(\mathbf{p}_{a},\mathbf{p}_{b})}\chi^{ab}_{S_{C}M_{S_{C}}}|q_{a}(\mathbf{p}_{a})q_{b}(\mathbf{p}_{b})\rangle.$ (3)
@@
×$\times\langle L_{A}M_{L_{A}};S_{A}M_{S_{A}}|J_{A},M_{J_{A}}\rangle\langle 1m;1-m|00\rangle.$
Considering the vertices given by the quark pair creation model too hard at high vertices, we use a momentum-dependent form factor to suppress the unphysical contributions as in the literatures [55, 56], and reads
@@
In the equation, $\Lambda=780$ MeV is the value used herein. p stands for the momentum of the daughter baryon $B$ in the center of mass frame of the parent baryon $A$, and reads
@@
$\Xi_{c}^{\prime+}$ 2578.2 $D^{0}$ 1864.84 $\Sigma_{c}|J^{P}=\frac{5}{2}^{-},2\rangle_{\lambda}$ 2935
@@
In this work, we adopt $m_{u}$=$m_{d}$=330 MeV, $m_{s}$=450 MeV, and $m_{c}$=1700 MeV for the constituent quark mass. The masses of the final baryons and mesons involved in our calculations, collected in Table 2. The harmonic oscillator strength $R=2.5$ GeV-1 for light flavor mesons $\pi/K^{(*)}/\omega/\eta$, $R=1.67$ GeV-1 for $D$ meson and $R=1.94$ GeV-1 for $D^{*}$ meson [58]. The parameter of the $\rho$-mode excitation between the two light quarks is taken as $\alpha_{\rho}$=0.4 GeV. The other parameter $\alpha_{\lambda}$ is obtained by the relation [44]
@@ III.1 $1P$-wave $\lambda$-mode excitations
Decay width $\Lambda_{c}\ket{J^{P}=\frac{1}{2}^{-},1}_{\lambda}$ $\Lambda_{c}\ket{J^{P}=\frac{3}{2}^{-},1}_{\lambda}$
@@ III.2 $1P$-wave $\rho$-mode excitations
For the $1P$-wave $\rho$-mode $\Lambda_{c}$ baryons, there are five states $\Lambda_{c}\ket{J^{P}=1/2^{-},0}_{\rho}$, $\Lambda_{c}\ket{J^{P}=1/2^{-},1}_{\rho}$, $\Lambda_{c}\ket{J^{P}=3/2^{-},1}_{\rho}$, $\Lambda_{c}\ket{J^{P}=3/2^{-},2}_{\rho}$ and $\Lambda_{c}\ket{J^{P}=5/2^{-},2}_{\rho}$. According to the theoretical predictions by various methods, the mass of the $1P$-wave $\rho$-mode $\Lambda_{c}$ baryons is about $M\sim$2.85 GeV. Meanwhile, we notice that their masses are above the threshold of $ND$, while their strong decays are forbidden due to the orthogonality of spatial wave functions. This is true for all of the $\rho$-mode excitations. Hence, we mainly focus on their strong decays into $\Sigma_{c}\pi$ and $\Sigma_{c}^{*}\pi$. Fixing the masses of $1P$-wave $\rho$-mode $\Lambda_{c}$ baryons at the predictions in Ref. [30], we study their strong decay properties, and list in Table 4.
@@
Decay width $\Lambda_{c}\ket{J^{P}=\frac{1}{2}^{-},0}_{\rho}$ $\Lambda_{c}\ket{J^{P}=\frac{1}{2}^{-},1}_{\rho}$ $\Lambda_{c}\ket{J^{P}=\frac{3}{2}^{-},1}_{\rho}$ $\Lambda_{c}\ket{J^{P}=\frac{3}{2}^{-},2}_{\rho}$ $\Lambda_{c}\ket{J^{P}=\frac{5}{2}^{-},2}_{\rho}$
@@
Decay width $\Lambda_{c1}\ket{J^{P}=\frac{1}{2}^{-},0}_{\rho}$ $\Lambda_{c1}\ket{J^{P}=\frac{1}{2}^{-},1}_{\rho}$ $\Lambda_{c1}\ket{J^{P}=\frac{3}{2}^{-},1}_{\rho}$ $\Lambda_{c1}\ket{J^{P}=\frac{3}{2}^{-},2}_{\rho}$ $\Lambda_{c1}\ket{J^{P}=\frac{5}{2}^{-},2}_{\rho}$
@@
$\Gamma[\Lambda_{c}\eta^{\prime}]$ 3.38 -
$\Gamma[\Sigma_{c}\rho]$ 0.23 0.11 176.79 -
@@
Within the $j-j$ coupling scheme, the total decay width of $\Lambda_{c}|J^{P}=1/2^{-},0\rangle_{\rho}$ is most likely to be near zero. We know that the states in the $j-j$ coupling scheme can be expressed with the linear combination of the configurations in the $L-S$ coupling scheme. That means the states in the $j-j$ coupling scheme will contain a mixing angle $\theta$. Considering the heavy quark symmetry being not strictly true and slightly breaking in the $\Lambda_{c}$ system, the mixing angle $\theta$ will fluctuate around the center value($\theta\simeq 35^{\circ}$). To investigate this effect, we plot the strong decay widths of $\Lambda_{c}|J^{P}=1/2^{-},0\rangle_{\rho}$ as a function of the mixing angle in Fig 2. we can obtained that $\Lambda_{c}|J^{P}=1/2^{-},0\rangle_{\rho}$ is still very narrow state, and the $\Sigma_{c}\pi$ decay channel almost saturates its total decay widths.
Figure 2: Partial and total strong decay widths of $\Sigma_{c}|J^{P}=1/2^{-},0\rangle_{\rho}$ as a function of the mixing angle.
@@
which is agreement with the observation as well. The dominant decay channel is $\Sigma_{c}^{*}\pi$, and the corresponding partial decay width ratio is
$\frac{\Gamma[\Lambda_{c}|J^{P}=\frac{5}{2}^{-},2\rangle_{\rho}]\rightarrow\Sigma_{c}\pi}{\Gamma[\Lambda_{c}|J^{P}=\frac{5}{2}^{-},2\rangle_{\rho}]\rightarrow\Sigma_{c}^{*}\pi}\simeq 0.51.$ (16)
Meanwhile, the predicted branching ratio the $\Sigma_{c}\pi$ channel is

### v2

@@ Abstract
Inspired by the latest experimental progress, we systematically study the Okubo-Zweig-Iizuka(OZI)-allowed two-body strong decay properties of $1P$-, $1D$-, $2S$- and $2P$-wave $\Lambda_{c}$ baryons within the $j$-$j$ coupling scheme in the framework of the quark pair creation model. The calculations indicate that: ($\rm{i}$) Taking the observed states $\Lambda_{c}(2595)^{+}$ and $\Lambda_{c}(2625)^{+}$ as the $1P$-wave $\lambda$-modes states $\Lambda_{c}|J^{P}=1/2^{-},1\rangle_{\lambda}$ and $\Lambda_{c}|J^{P}=3/2^{-},1\rangle_{\lambda}$, respectively, we can reproduce the experimental data well in theory. ($\rm{ii}$) Combining with the measured mass and the decay properties of $\Lambda_{c}(2860)^{+}$, this excited state can be explained as $1D$-wave $\lambda$-mode state $\Lambda_{c}|J^{P}=3/2^{+},1\rangle_{\lambda\lambda}$. ($\rm{iii}$) The newly observed state $\Lambda_{c}(2910)^{+}$ may be assigned as one of the $1P$-wave $\rho$-mode states $\Lambda_{c}|J^{P}=3/2^{-},2\rangle_{\rho}$ or $\Lambda_{c}|J^{P}=5/2^{-},2\rangle_{\rho}$. Meanwhile, we notice that the partial decay width ratio between $\Sigma_{c}\pi$ and $\Sigma_{c}^{*}\pi$ for the two candidates is significantly different. Hence, experimental progress in this ratio measurement may shed light on the nature of $\Lambda_{c}(2910)^{+}$. ($\rm{iv}$) According to the properties of $\Lambda_{c}(2765)^{+}$, we find that the $2S$-wave $\lambda$-mode state $\Lambda_{c1}|J^{P}=1/2^{+},0\rangle_{\lambda}$ may be a potential candidate. ($\rm{v}$) The $2P$-wave $\lambda$-mode state $\Lambda_{c1}|J^{P}=3/2^{-},1\rangle_{\lambda}$ is mostly likely to be a good assignment of the controversial state $\Lambda_{c}(2940)^{+}$. Both the total decay width and partial decay ratio between $pD^{0}$ and $\Sigma_{c}\pi$ are in good agreement with the observations. ($\rm{vi}$) In addition, for the missing $\Lambda_{c}$ excitations, we obtain their strong decay properties and hope that’s useful for future experimental exploration.
@@ I Introduction
Meanwhile, there exist many theoretical calculations to decode the inner structures of those observed $\Lambda_{c}^{+}$ baryons via mass spectrum [18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33] and decay properties [34, 33, 35, 36, 37, 38, 39, 19, 40, 41, 42, 43, 44, 45, 46, 47]. Except for the $\Lambda_{c}^{+}$ ground state, the $\Lambda_{c}(2595)^{+}$ and $\Lambda_{c}(2625)^{+}$ can be well interpreted as the $P$-wave $\Lambda_{c}$ states of $J^{P}=1/2^{-}$ and $J^{P}=3/2^{-}$ [18, 19, 20, 34, 35, 36, 37, 38, 39], respectively. Meanwhile, for the $\Lambda_{c}(2860)^{+}$ and $\Lambda_{c}(2880)^{+}$, most of the references suggest they together form the $D$-wave $\Lambda_{c}$ doublet of $J^{P}=3/2^{+}$ and $J^{P}=5/2^{+}$, respectively [25, 23, 26, 27, 41, 42]. However, there is some literature that don’t support $\Lambda_{c}(2880)^{+}$ as $J^{P}=5/2^{+}$ state [27, 41], and believe this state being $F$-wave $\Lambda_{c}$ states of $J^{P}=5/2^{-}$ [41]. Hence, the relation between those two states still needs to be carefully examined in future experimental and theoretical studies. As for $\Lambda_{c}(2940)^{+}$, its internal structure is controversial. In addition to being interpreted as a traditional hadron state of $J^{P}=1/2^{-}$ [2], $3/2^{\pm}$ [22, 24, 41, 43], $5/2^{\mp}$ [43, 45] or $7/2^{+}$ [44], it has also been interpreted as $D^{*}N$ molecular state with $J^{P}=3/2^{-}$ [48, 49]. Meanwhile, there are some references [50, 51] discussing the properties of $\Lambda_{c}(2940)$ for helping us to clarify its nature. Fortunately, the latest experimental measurements by Belle Collaboration [16] about the partial decay width ratios of $\Lambda_{c}(2940)^{+}$ will provide a stronger basis for decoding its inner structure. Compared to $\Lambda_{c}(2940)^{+}$, the properties of $\Lambda_{c}(2765)^{+}$(or $\Sigma_{c}(2765)^{+}$) and $\Lambda_{c}(2910)^{+}$ are more controversial, and it is not even certain whether they are $\Lambda_{c}$ or $\Sigma_{c}$ states. At present, theoretical explanations suggest that $\Lambda_{c}(2765)^{+}$(or $\Sigma_{c}(2765)^{+}$) may be $\Lambda_{c}(2S)1/2^{+}$ [21, 22], $\Lambda_{c}(1P)1/2^{-}$ [22] or $\Sigma_{c}(1P)3/2^{-}$ [21] resonance. For $\Lambda_{c}(2910)^{+}$, it can be assigned as $\Lambda_{c}$ resonance with spin-parity $1/2^{+}$ [22] or $1/2^{-}$ [24]. Moreover, the assignment of the two states as $D^{(*)}N$ molecular states also exists [49, 52].
Table 1: Predicted masses of $\Lambda_{c}$ states($1P$-, $1D$-, $2S$- and $2P$-wave) in various quark models and quark pair creation modes.
@@
[…]
@@ II Theoretical framework
The quark pair creation model [53, 54, 55, 56, 57] as a phenomenological method has been employed successfully in the description of the OZI-allowed two-body strong decays. The main idea of this model is that the quark-antiquark pair with $0^{++}$ is created from the vacuum and then regroups with the quarks from the initial hadron to produce two outing hadrons. Hence, for the $\Lambda_{c}$ system, there are three decay processes, as shown in Fig 1.
@@
[…]
$\times\omega^{45}_{0}\chi^{45}_{1,-m}\phi^{45}_{0}\mathcal{Y}^{m}_{1}(\dfrac{\textbf{p}_{4}-\textbf{p}_{5}}{2})a^{\mathcal{y}}_{4i}(\textbf{p}_{4})b^{\mathcal{y}}_{5j}(\textbf{p}_{5}).$
@@
$\times\int d^{3}\mathbf{p}_{1}d^{3}\mathbf{p}_{2}d^{3}\mathbf{p}_{3}\delta^{3}(\mathbf{p}_{1}+\mathbf{p}_{2}+\mathbf{p}_{3}-\mathbf{p}_{A})$
$\times\Psi_{N_{A}L_{A}M_{L_{A}}(\mathbf{p}_{1},\mathbf{p}_{2},\mathbf{p}_{3})}\chi^{123}_{S_{A}M_{S_{A}}}|q_{1}(\mathbf{p}_{1})q_{2}(\mathbf{p}_{2})q_{3}(\mathbf{p}_{3})\rangle,$ (2)
@@
$\times\int d^{3}\mathbf{p}_{a}d^{3}\mathbf{p}_{b}\delta^{3}(\mathbf{p}_{a}+\mathbf{p}_{b}-\mathbf{p}_{C})$
$\times\Psi_{N_{C}L_{C}M_{L_{C}}(\mathbf{p}_{a},\mathbf{p}_{b})}\chi^{ab}_{S_{C}M_{S_{C}}}|q_{a}(\mathbf{p}_{a})q_{b}(\mathbf{p}_{b})\rangle.$ (3)
@@
[…]
Considering the vertex given by the quark pair creation model too hard at high momenta, we modify the vertex by adopting a form factor $e^{-\frac{\mathbf{p}^{2}}{2\Lambda^{2}}}$ as in the literatures [59, 60, 61], which gives the quark-pair-creation vertex a finite-size rather than point-like behavior. It reads
@@
In the equation, we fix the cut-off parameter $\Lambda=780$ MeV, which is the same as to the value used in Ref. [62]. p stands for the momentum of the daughter baryon $B$ in the center of mass frame of the parent baryon $A$, and reads
@@
$\Xi_{c}^{\prime+}$ 2578.2 $D^{0}$ 1864.84 $\Sigma_{c}|J^{P}=\frac{5}{2}^{-},2\rangle_{\lambda}$ 2835
@@
In this work, we adopt $m_{u}$$m_{d}$=330 MeV, $m_{s}$=450 MeV, and $m_{c}$=1700 MeV for the constituent quark mass. The masses of the final baryons and mesons involved in our calculations, collected in Table 2. The harmonic oscillator strength $R=2.5$ GeV-1 for light flavor mesons $\pi/K^{(*)}/\omega/\eta$, $R=1.67$ GeV-1 for $D$ meson and $R=1.94$ GeV-1 for $D^{*}$ meson [64]. The parameter of the $\rho$-mode excitation between the two light quarks is taken as $\alpha_{\rho}$=0.4 GeV. The other parameter $\alpha_{\lambda}$ is obtained by the relation [45]
@@ III.1 $1P$-wave $\lambda$-mode excitations
[…]
@@ III.2 $1P$-wave $\rho$-mode excitations
[…]
@@
[…]
@@
[…]
@@
$\Gamma[\Lambda_{c}\eta^{\prime}]$ - - - 3.38 -
$\Gamma[\Sigma_{c}\rho]$ - 0.23 0.11 176.79 -
@@
Within the jjj-j coupling scheme, the total decay width of $\Lambda_{c}|J^{P}=1/2^{-},0\rangle_{\rho}$ is most likely to be near zero. We know that the states in the jjj-j coupling scheme can be expressed with the linear combination of the configurations in the LSL-S coupling scheme, which reads
$\left|\left\{[(l_{\rho}l_{\lambda})_{L}s_{\rho}]_{j}s_{Q}\right\}_{J^{P}}\right\rangle$ $=(-1)^{L+s_{\rho}+\frac{1}{2}+J}\sqrt{2J+1}\sum_{S}\sqrt{2S+1}$ (13)
$\begin{pmatrix}L&s_{\rho}&j\\ s_{Q}&J&S\end{pmatrix}\left|\left\{[(l_{\rho}l_{\lambda})_{L}(s_{\rho}s_{Q})_{S}]_{J}\right\}\right\rangle.$
In the expression, $l_{\rho}$ and $l_{\lambda}$ are the $\rho$- and $\lambda$-modes quantum numbers of the orbital angular, respectively. The total orbital angular momentum $L$ = $|l_{\rho}-l_{\lambda}|,\cdots,l_{\rho}+l_{\lambda}$. $s_{\rho}$ is the quantum numbers of the total spin of the two light quarks and $s_{Q}$ is the spin of the heavy quark. The total spin angular momentum $S$ = $|s_{\rho}-s_{Q}|,\cdots,s_{\rho}+s_{Q}$. $J$ is the total angular momentum. That means the states in the jjj-j coupling scheme will contain a mixing angle $\theta$. Considering the heavy quark symmetry being not strictly true and slightly breaking in the $\Lambda_{c}$ system, the mixing angle $\theta$ will fluctuate around the center value($\theta\simeq 35^{\circ}$). To investigate this effect, we plot the strong decay widths of $\Lambda_{c}|J^{P}=1/2^{-},0\rangle_{\rho}$ as a function of the mixing angle in Fig 2. we can obtained that $\Lambda_{c}|J^{P}=1/2^{-},0\rangle_{\rho}$ is still very narrow state, and the $\Sigma_{c}\pi$ decay channel almost saturates its total decay widths.
Figure 2: Partial and total strong decay widths of $\Lambda_{c}|J^{P}=1/2^{-},0\rangle_{\rho}$ as a function of the mixing angle.
@@
which is agreement with the observation as well. The dominant decay channel is $\Sigma_{c}^{*}\pi$, and the corresponding branching fraction is
$\frac{\Gamma[\Lambda_{c}|J^{P}=\frac{5}{2}^{-},2\rangle_{\rho}]\rightarrow\Sigma_{c}^{*}\pi}{\Gamma_{\mathrm{Total}}}\simeq 66\%.$ (17)
Meanwhile, the predicted branching ratio of the $\Sigma_{c}\pi$ channel is
@@
Of course, considering that the current experimental data are limited and bare large errors, there are other possible explanations, such as $D^{*}N$ molecular state [49, 52] or $\Lambda_{c}$ resonances with different spin-parity [22, 24]. To further clarify the nature of $\Lambda_{c}(2910)^{+}$, more discussions are necessary.